### v1

# thanks: Corresponding author

Abstract
In the framework of mutually unbiased bases (MUBs), a measurement in one basis gives no information about the outcomes of measurements in another basis. Here, we relax the no-information condition by allowing the $d$ outcomes to be predicted according to a predefined probability distribution $q=(q_{0},\cdots,q_{d-1})$. The notion of mutual unbiasedness, however, is preserved by requiring that the extracted information is the same for any preparation and any measurement; regardless of which state from which basis is chosen to prepare the system, the outcomes of measuring the system with respect to the other basis generate the same probability distribution. In the light of this, we define the notion of mutually equi-biased bases (MEBs) such that within each basis the states are equi-biased with respect to the states of the other basis and that the bases are mutually equi-biased with respect to each other. For $d=2,3$, we derive a complete set of $d+1$ MEBs. The mutual equi-biasedness imposes nontrivial constraints on the distribution $q$, leading for $d=3$ to the restriction $1/3\leq\mu\leq 1/2$ where $\mu=\sum_{k=0}^{2}q_{k}^{2}$. To capture the incompatibility of the measurements in MEBs, we derive an inequality for the probabilities of projective measurements in a qudit system, which yields an associated entropic uncertainty inequality. Finally, we construct a class of positive maps and their associated entanglement witnesses based on MEBs. While an entanglement witness constructed from MUBs is generally finer than one based on MEBs when both use the same number of bases, for certain values of the index $\mu$, employing a larger set of MEBs can yield a finer witness. We illustrate this behavior using isotropic states of a $3\times 3$ system.
mutually unbiased bases; mutually equi-biased bases; entropic uncertainty relation; entanglement witness
I Introduction
Quantum mechanics is full of mysteries, and perhaps the most profound is the existence of incompatible measurements, which imposes fundamental limitations on the precision of measurement outcomes through uncertainty and complementarity HeisenbergZfP1927 ; BohrNature1928 . This restricts the amount of information that can be extracted from measurements on a system, and it makes sense, therefore, that the more incompatible two measurements are, the greater the restrictions on the precision with which certain properties of a system can be simultaneously known. The uncertainty relation quantifies how this incompatibility restricts the precision with which we can simultaneously measure incompatible observables. Related to this, mutually unbiased bases (MUBs) play a crucial role, as measurements in one basis provide no information about the outcomes of measurements in another basis SchwingerPNAS1960 ; WoottersAP1989 ; IvanovicJPA1981 .
[…]
${\textrm{Tr}}[P_{k}^{(\alpha)}P_{l}^{(\beta)}]=1/d+\delta_{\alpha\beta}(\delta_{kl}-1/d),$ (1)
for $k,l=0,\cdots,d-1$ and $\alpha,\beta=1,\cdots,L$. Accordingly, if a system is prepared in a state belonging to one of the bases, then the outcomes of the measuring the system with respect to the other basis give, by no means, any information about the state of the system, i.e., all outcomes are predicted to occur with equal probability.
For a $d$-dimensional system, the number of MUBs is at most $d+1$, accordingly $L\leq d+1$. However, except for the particular cases where the dimension $d$ is a prime or a power of a prime number, for which there exist a complete set of $d+1$ MUBs, the maximum number of MUBs in other dimensions is not known. For a general dimension $d$, it is known that at least three MUBs always exist. It is shown that a complete set of MUBs, if exists, forms a minimal and optimal set of projective measurements for quantum state determination WoottersAP1989 ; IvanovicJPA1981 . The mutual unbiasedness implies that the statistical errors are minimized when finite samples are measured WoottersAP1989 ; ZyczkowskiIJQI2010 ; RenesJMP2004 .
By generalizing the notion of unbiasedness from projective measurements to more general positive operator-valued measures (POVMs), Kalev et al. KalevNJP2014 introduced the notion of mutually unbiased measurements (MUMs), which encompass a broader class of measurements that retain the essential features of unbiasedness. Mutually unbiased bases and measurements have found important applications in quantum information, serving both as effective methods for entanglement detection SpenglerPRA2012 ; ChruscinskiPRA2014 ; ChenPRA2014 ; ChruscinskiPRA2018 ; LiIJTP2019 ; SalehiQIP2021 , and as foundational frameworks for entropic certainty and uncertainty relations MaassenPRL1988 ; ColesRMP2017 ; BallesterPRA2007 ; MolmerPRA2009 ; WehnerNJP2010 ; SanchezPLA1993 ; ZyczkowskiPRA2015 . Whereas entropic uncertainty relations set lower bounds on average entropies, reflecting the impossibility of a state being a simultaneous eigenstate of all observables, entropic certainty relations set upper bounds, arising from the fact that—apart from the completely mixed state—no quantum state can produce $d$ equally likely outcomes for every observable at once SanchezPLA1993 . For the average entropy characterizing measurements of a pure quantum state of dimension $d$ in $L$ orthogonal bases, both lower and upper bounds have been established in ZyczkowskiPRA2015 .
Here we extend this notion of MUBs to define mutually equi-biased bases (MEBs) such that the standard MUBs are obtained as a special limiting case of our results. To this aim, the unbiasedness condition is partially relaxed for MEBs in the sense that the transition probability from any state of one basis to any state of the second basis is no longer necessarily equal to $1/d$. In the most general setting of quantum measurement, if a system is prepared in a state belonging to one basis, the statistics of the outcomes of the measuring the system with respect to the states of the other basis provides a probability distribution ${q}=(q_{0},\cdots,q_{d-1})$. We adapt this notion of biasedness, however, our restriction is that regardless of which state of the first basis is chosen, the probability distribution of the transition to each state of the second basis remains the same up to a permutation of its entries. Accordingly, within each basis the states are equi-biased with respect to the states of the other basis and that the bases are mutually equi-biased with respect to each other. In the light of this, we associate a probability distribution ${q}=(q_{0},\cdots,q_{d-1})$ with each of MEBs such that
${\textrm{Tr}}[P_{k}^{(\alpha)}P_{l}^{(\beta)}]=q_{\sigma^{(}k)\oplus\sigma(l)}+\delta_{\alpha\beta}(\delta_{kl}-q_{\sigma(k)\oplus\sigma(l)}),$ (2)
which reduces to Eq. (1) of MUBs when $q_{k}=1/d$ for $k=0,\cdots,d-1$. Here, $\oplus$ stands for the sum module $d$, and $\sigma(k)$ denotes the permutation of $k\in(0,1,\cdots,d-1)$.
In MEBs, in contrary to MUBs, the knowledge of the measuring the system with respect to the one basis, when the system is prepared in any state of the other basis, leaks some information about the state of the system. The mutual equi-biasedness of the bases implies, however, that they are biased in the same degree in the sense that regardless of which state from which basis are chosen to prepare the system, the extracted information is the same, i.e., they are predictable in the same degree with respect to each other. It makes sense therefore that the question of how much information can be extracted should be related to how far the probability distribution is from a homogenous distribution. The index of coincidence, defined by $\mu(q)=\sum_{k=0}^{d-1}q_{k}^{2}$ for the probability $q=(q_{0},\cdots,q_{d-1})$, is a measure of how much a probability distribution deviates from a uniform distribution, as such, it can play a key role when the extracted information is quantified. Hereafter, for brevity, we will refer to the index of coincidence simply as the index and denote it by $\mu$ for a $d$-outcomes probability distribution $q=(q_{0},\cdots,q_{d-1})$.
Related to the index $\mu$, we now pose the question that whether for a given $d$, a complete set of $d+1$ MEBs exists for all values of $\mu\in[1/d,1]$? We derive a complete set of MEBs for two cases $d=2,3$. Our results show that except for the particular case of $2$-dimensional system for which a complete set of $3$ MEBs exists for all $\mu\in[1/2,1]$, it is not the case for $d\geq 3$. More precisely, for $d=3$, the constraints imposed by the mutual equi-biasedness limits the range of the index $\mu$ to the feasible region $1/3\leq\mu\leq 1/2$ within which a set of four MEBs exists.
Quantum incompatibility imposes limitation on the information that can be extracted from measurements of different observables, and the entropic uncertainty relation provides a bound on how much information can be gained from such measurements. To capture this notion of incompatibility in MEBs, we derive an inequality for the probabilities of projective measurements in a qudit system, from which we obtain an entropic uncertainty inequality.
Among the various applications of MUBs, their role in entanglement detection is particularly important. Similarly, MEBs can also be used for identifying entanglement. The most general tool to detect entanglement is the so-called entanglement witness, and a powerful tool to construct entanglement witness is to use positive but not completely positive maps. Based on MEBs, we construct a class of positive maps and their associated entanglement witnesses.
The paper is organized as follows. In Sec. II, we introduce the notion of MEBs and present a complete derivation of MEBs for the two simplest cases, $d=2,3$. In Sec. III, we derive an inequality for the probabilities of projective measurements in MEBs of a qudit system. Based on this, we provide an entropic uncertainty inequality in MEBs. In Sec. IV, we construct a positive map based on MEBs, then introduce an entanglement witness and examine its performance for two-qutrit isotropic states. Finally, we conclude the paper in Sec. V with a brief summary. Some proofs of the results are provided in Appendices A to C.
II Mutually equi-biased bases (MEBs)
II.1 Preliminaries
We begin this section by defining the notion of mutually equi-biased bases (MEBs), which generalizes the concept of mutually unbiased bases (MUBs) to encompass a broader family of probability distributions. We will see that the allowed region of this family is shaped by the mutually equi-biased condition.
Definition 1.
Let $q=(q_{0},\cdots,q_{d-1})$ be a probability distribution over $d$ outcomes. We say that the set of $L$ orthonormal bases $\{\mathcal{B}^{(1)},\cdots,\mathcal{B}^{(L)}\}$ of the $d$-dimensional Hilbert space $\mathcal{H}$ are MEBs associated to the probability distribution $q$, if for every pair $\alpha\neq\beta$, the corresponding projection operators satisfy Eq. (2) for some permutation $\sigma$. Here, $\mathcal{B}^{(\alpha)}=\{|e_{k}^{(\alpha)}\rangle\}$ where $\alpha=1,\cdots,L$, and $k=0,\cdots,d-1$.
The definition of MEBs relies on a probability distribution $q=(q_{0},\cdots,q_{d-1})$, from which one can define the $d\times d$ matrix $\mathcal{Q}$ with entries $\mathcal{Q}_{kl}=q_{k\oplus l}$. Clearly, $\mathcal{Q}$ is bistochastic; its entries are nonnegative with each of its rows and columns summing to unity, i.e.,
$\mathcal{Q}_{kl}\geq 0,\qquad\sum_{k=0}^{d-1}\mathcal{Q}_{kl}=1,\qquad\sum_{l=0}^{d-1}\mathcal{Q}_{kl}=1.$ (3)
More precisely, the set of matrices $\mathcal{Q}$s forms a (d1)(d-1)-dimensional subclass of the bistochastic matrices of size $d\times d$, which themselves comprise a convex set of dimension $(d-1)^{2}$, known as Birkhoff’s polytope ZyczkowskiCMP2005 . A bistochastic matrix $\mathcal{Q}$ is called unistochastic matrix if and only if its entries are the squares of the absolute values of the entries of some unitary matrix. In addition to the normalization requirement—i.e., in each row and column of a unitary matrix, the squared moduli sum to unity—the moduli of its entries must also satisfy certain constraints arising from the mutual orthogonality of the rows and columns. Accordingly, not all bistochastic matrices arise from unitary matrices; nonetheless, the $d\times d$ bistochastic matrix $\mathcal{Q}$, associated with the transition probabilities between two orthonormal bases $\mathcal{B}^{(\alpha)}$ and $\mathcal{B}^{(\beta)}$ is a unistochastic matrix. The characterization of the unistochastic subset of Birkhoff’s polytope, i.e., those for which there exist a unitary matrix $\mathcal{U}$ such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}|^{2}$, is considered in ZyczkowskiCMP2005 .
To set the convention, suppose $\mathcal{B}^{(1)}=\{|e_{k}^{(1)}\rangle\}_{k=0}^{d-1}$ denotes the standard (computational) basis of the $d$-dimensional Hilbert space $\mathcal{H}$. Then, for each $\beta\in\{2,\cdots,L\}$, the orthonormal basis $\mathcal{B}^{(\beta)}$ is related to $\mathcal{B}^{(1)}$ via a unitary matrix $\mathcal{U}^{(\beta)}$ as $|e_{k}^{(\beta)}\rangle=\mathcal{U}^{(\beta)}|e_{k}^{(1)}\rangle$, for $k=0,\cdots,d-1$, where $\mathcal{U}^{(\beta)}_{kl}=\langle e_{k}^{(1)}|e_{l}^{(\beta)}\rangle$. By equi-biasedness, we have to assign $|\mathcal{U}_{kl}^{(\beta)}|^{2}=\mathcal{Q}_{kl}$, for all $\beta\in\{2,\cdots,L\}$, where the $d\times d$ bistochastic matrix $\mathcal{Q}$ is defined by $\mathcal{Q}_{kl}=q_{k\oplus l}$ for the probability distribution $q=(q_{0},\cdots,q_{d-1})$. An important question is to determine for which probability distributions $q$ is the associated bistochastic matrix $\mathcal{Q}$ unistochastic—that is, which $q$ gives rise to unistochastic matrix $\mathcal{Q}$ corresponding to unitary matrices $\mathcal{U}^{(\beta)}$ such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}^{(\beta)}|^{2}$?
Before we proceed further, let us first note that for a given unistochastic matrix $\mathcal{Q}$, if one unitary matrix $\mathcal{U}$ is found such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}|^{2}$, then there exist infinitely many unitary matrices that satisfy this condition. This follows from the fact that the left and right multiplication of a unitary matrix $\mathcal{U}$ by any diagonal unitary matrices $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ results in another unitary matrix $\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$, but the associated unistochastic matrix remains unchanged ZyczkowskiJMP2009 . What is crucial in our analysis, however, is the requirement of the mutual equi-biasedness. This condition restricts the set of MEBs to be finite. To be more precise, suppose $\mathcal{Q}$ is a unistochastic matrix generated by L1L-1 unitary matrices $\mathcal{U}^{(\beta)}$, such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}^{(\beta)}|^{2}$, for each $\beta\in\{2,\cdots,L\}$. The mutually equi-biasedness requires further that they are mutually unistochastic, i.e., $|[{\mathcal{U}^{(\alpha)}}^{\dagger}\mathcal{U}^{(\beta)}]_{kl}|^{2}=\mathcal{Q}_{\sigma(k)\sigma(l)}$ for all $\alpha,\beta\in\{2,\cdots,L\}$ with $\alpha\neq\beta$. The permutation $\sigma$, appearing here and in Eq. (2), depends on $\alpha$ and $\beta$ in general, however, when one of the basis corresponds to the standard basis, we follow the convention that the permutation $\sigma$ does nothing, i.e., $\sigma(k)=k$ and $\sigma(l)=l$.
In the following subsections, we examine two special cases of $d=2$ and $d=3$, providing explicite construction of MEBs in each case.
II.2 Mutually equi-biased bases for $d=2$
For the simplest case of two-level system, suppose $\mathcal{B}^{(1)}=\{|0\rangle,|1\rangle\}$ represents the standard basis. Associated with the probability distribution ${q}=(q_{0},q_{1})$, or its corresponding bistochastic matrix $\mathcal{Q}=\begin{pmatrix}q_{0}&q_{1}\\ q_{1}&q_{0}\end{pmatrix}$, we have to construct the unitary matrix $\mathcal{U}$ with prescribed moduli of its matrix elements. Starting with $\mathcal{U}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}\\ \sqrt{q_{1}}&\sqrt{q_{0}}{\mathrm{e}}^{i\xi}\end{pmatrix}$, unitarity requires $\xi=\pi$, hence
$\mathcal{U}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}\\ \sqrt{q_{1}}&-\sqrt{q_{0}}\\ \end{pmatrix}$. (4)
Given that for any diagonal unitary matrices $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, the unistochastic matrix associated with $\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$ is the same as that associated with $\mathcal{U}$ ZyczkowskiJMP2009 , one can construct a second unitary matrix $\mathcal{V}$ from $\mathcal{U}$ as $\mathcal{V}=D\mathcal{U}$ where $D={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha}\}$. We obtain
$\mathcal{V}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}\\ \sqrt{q_{1}}{\mathrm{e}}^{i\alpha}&-\sqrt{q_{0}}{\mathrm{e}}^{i\alpha}\\ \end{pmatrix}$. (5)
Although both $\mathcal{U}$ and $\mathcal{V}$ generate the same bistochastic matrix $\mathcal{Q}$ and both define bases that are MEBs with respect to the standard basis, they are not necessarily MEBs with respect to each other. For this to hold, $\mathcal{U}$ and $\mathcal{V}$ must be mutually equi-biased, i.e., the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$ must generate the same bistochastic matrix $\mathcal{Q}$. This places a constraint on $\alpha$ such that $q_{0}^{2}+q_{1}^{2}+2q_{0}q_{1}\cos{\alpha}=q_{0}$, where uniquely determines $\alpha$ as
$\alpha=\cos^{-1}[\delta/(1+\delta)].$ (6)
Here, $\delta=q_{0}-q_{1}$ is introduced both as an independent parameter for the binary distribution $q=(q_{0},q_{1})$ (see Appendix C for the underlying motivation) and as a measure of its unbiasedness.
Starting from the standard orthonormal basis $\{|e_{k}^{(1)}\rangle=|k\rangle\}$, one can apply the unitary matrices $\mathcal{U}$ and $\mathcal{V}$ to construct new bases $\{|e_{k}^{(2)}\rangle=\mathcal{U}|e_{k}^{(1)}\rangle\}$ and $\{|e_{k}^{(3)}\rangle=\mathcal{V}|e_{k}^{(1)}\rangle\}$, for $k=0,1$. This notion of equi-biasedness admits an interesting geometric interpretation in the Bloch vector representation, in the sense that with $P_{k}^{(\alpha)}=({\mathbb{I}}_{2}+\boldsymbol{r}_{k}^{(\alpha)}\cdot\boldsymbol{\sigma})/2$ as the Bloch vector representation of the projection operator $P_{k}^{(\alpha)}$, the Bloch vectors satisfy
$\boldsymbol{r}_{k}^{(\alpha)}\cdot\boldsymbol{r}_{l}^{(\beta)}=(-1)^{k\oplus l}\;\delta,$ $\mbox{if }\;\alpha\neq\beta.$ (7)
Accordingly, the Bloch vectors form a symmetric or equiangular triple such that they are mutually symmetric with the same mutual angle $\delta$ (see Fig. 1).
Figure 1: Bloch vector representation for MEB, when $d=2$. Geometrically, equi-biasedness means equal angle between Bloch vectors of different bases.
II.3 Mutually equi-biased bases for $d=3$
In this case, associated with the probability distribution $q=(q_{0},q_{1},q_{2})$, we have to construct the unitary matrix $\mathcal{U}$ with prescribed moduli of its matrix elements as
$\mathcal{U}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}&\sqrt{q_{2}}\\ \sqrt{q_{1}}&\sqrt{q_{2}}{\mathrm{e}}^{i\xi_{0}}&\sqrt{q_{0}}{\mathrm{e}}^{i\xi_{1}}\\ \sqrt{q_{2}}&\sqrt{q_{0}}{\mathrm{e}}^{i\eta_{1}}&\sqrt{q_{1}}{\mathrm{e}}^{i\eta_{0}}\end{pmatrix}$. (8)
Mutual orthogonality of the rows (and columns) gives simply the following conditions
$L_{2}+L_{0}{\mathrm{e}}^{i\xi_{0}}+L_{1}{\mathrm{e}}^{i\xi_{1}}=0,$ (9)
$L_{1}+L_{2}{\mathrm{e}}^{i\eta_{1}}+L_{0}{\mathrm{e}}^{i\eta_{0}}=0,$ (10)
$L_{0}+L_{1}{\mathrm{e}}^{i(\xi_{0}-\eta_{1})}+L_{2}{\mathrm{e}}^{i(\xi_{1}-\eta_{0})}=0,$ (11)
where we have defined ZyczkowskiJMP2009
$L_{0}=\sqrt{q_{1}q_{2}},\quad L_{1}=\sqrt{q_{2}q_{0}},\quad L_{2}=\sqrt{q_{0}q_{1}}.$ (12)
Equations (9) to (11) imply that the unitarity is fulfilled if we find the phases in such a way that three line segments $L_{0},L_{1},L_{2}$ form a triangle. This imposes the triangle inequality on the line segments ZyczkowskiCMP2005
$|L_{j}-L_{k}|\leq L_{i}\leq L_{j}+L_{k},$ (13)
where $(i,j,k)$ are different choices of $(0,1,2)$. For probabilities $q=(q_{0},q_{1},q_{2})$ for which the above inequalities hold, Eqs. (9) to (11) give the phases as
$\cos{\xi_{0}}$ $=$ $\frac{L_{1}^{2}-L_{0}^{2}-L_{2}^{2}}{2L_{0}L_{2}}=\cos{(\eta_{0}-\eta_{1})},$ (14)
$\cos{\xi_{1}}$ $=$ $\frac{L_{0}^{2}-L_{1}^{2}-L_{2}^{2}}{2L_{1}L_{2}}=\cos{\eta_{1}},$ (15)
$\cos{(\xi_{0}-\xi_{1})}$ $=$ $\frac{L_{2}^{2}-L_{0}^{2}-L_{1}^{2}}{2L_{0}L_{1}}=\cos{\eta_{0}}.$ (16)
A subset of the triangle probability that satisfies the chain-link conditions (13)—so that the matrix $\mathcal{Q}$ is unistochastic and can be generated from a unitary matrix—is shown in Fig. 2(a). The figure is plotted in the $\delta_{2}\delta_{1}$-plane, where the two independent parameters $\delta_{1}$ and $\delta_{2}$ are defined as follows (see Appendix C)
$q_{0}$ $=$ $1/3+2\delta_{1}/3,$ (17)
$q_{1}$ $=$ $1/3-\delta_{1}/3-\delta_{2}/\sqrt{3},$
$q_{2}$ $=$ $1/3-\delta_{1}/3+\delta_{2}/\sqrt{3}.$
Equivalently
$\delta_{1}=(3q_{0}-1)/2,\qquad\delta_{2}=\sqrt{3}(q_{2}-q_{1})/2,$ (18)
such that
$-1/2\leq$ $\delta_{1}$ $\leq 1,$ (19)
$(-1+\delta_{1})/\sqrt{3}\leq$ $\delta_{2}$ $\leq(1-\delta_{1})/\sqrt{3}.$ (20)
Figure 2: Geometric representation of the probability distribution $q=(q_{0},q_{1},q_{2})$ in the $\delta_{2}\delta_{1}$-plane. (a) A subset of the triangle probability where the bistochastic matrix $\mathcal{Q}$ is a unistochastic, i.e., the probabilities for which the conditions (13) are satisfied. (b) The feasible region that satisfy the constraint $\mu\leq 5/9$. (c) The feasible domain, after imposing the chain-linked conditions (26). (d) The portion of the feasible region for which $\mu\leq 1/2$.
One can construct the second unitary matrix $\mathcal{V}$ with the same prescribed moduli of its matrix elements as $\mathcal{U}$ as $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$ where $\mathcal{D}_{\alpha}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha_{1}},{\mathrm{e}}^{i\alpha_{2}}\}$. Regarding the probability $q=(q_{0},q_{1},q_{2})$, although both $\mathcal{U}$ and $\mathcal{V}$ generate bases that are equi-biased with respect to the standard basis, they are not necessarily equi-biased with respect to each other. For this to hold, $\mathcal{U}$ and $\mathcal{V}$ must be mutually equi-biased, i.e., the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$ must generate the same probability matrix $\mathcal{Q}$, up to a permutation of its entries. This places some constraints on phases $\alpha_{1}$ and $\alpha_{2}$. Furthermore, the mutual equi-biasedness imposes some constraints on the index $\mu=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}$ of the distribution $q=(q_{0},q_{1},q_{2})$—Not any value of $\mu$ is allowed to generate a bistochastic matrix from unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$.
To proceed with this claim, we have to compute the matrix $\mathcal{V}^{\dagger}\mathcal{U}$ and require that the squared moduli of its diagonal entries generate the probability distribution $q^{\sigma}=(q_{\sigma(0)},q_{\sigma(1)},q_{\sigma(2)})$, such that
$|[\mathcal{V}^{\dagger}\mathcal{U}]_{00}|^{2}=q_{\sigma(0)},\;|[\mathcal{V}^{\dagger}\mathcal{U}]_{11}|^{2}=q_{\sigma(2)},\;|[\mathcal{V}^{\dagger}\mathcal{U}]_{22}|^{2}=q_{\sigma(1)}.$ (21)
Here, $\sigma\in S_{3}$ denotes a permutation of three elements $\{0,1,2\}$, where $S_{3}$ is the symmetric group of degree $3$. For further use, recall that the $3!$ group operations of $S_{3}$ can be denoted as: the identity $\mathrm{id}$, the pair permutations $(12)$, $(13)$, $(23)$, and the forward and backward cycles $(123)$ and $(321)$. After some calculations, we get
$\mu+2q_{0}q_{1}\cos{\alpha_{1}}+2q_{0}q_{2}\cos{\alpha_{2}}+2q_{1}q_{2}\cos{\alpha_{12}}=q_{\sigma(0)},$ (22)
$\mu+2q_{1}q_{2}\cos{\alpha_{1}}+2q_{0}q_{1}\cos{\alpha_{2}}+2q_{0}q_{2}\cos{\alpha_{12}}=q_{\sigma(2)},$ (23)
$\mu+2q_{0}q_{2}\cos{\alpha_{1}}+2q_{1}q_{2}\cos{\alpha_{2}}+2q_{0}q_{1}\cos{\alpha_{12}}=q_{\sigma(1)},$ (24)
where we have defined $\alpha_{12}=\alpha_{1}-\alpha_{2}$, for the sake of brevity. Summing the three equations above, we obtain
$\cos{\alpha_{1}}+\cos{\alpha_{2}}+\cos{\alpha_{12}}=\frac{1-3\mu}{1-\mu}.$ (25)
For $\mu\in[1/3,1)$, the rhs of Eq. (25) can take any nonpositive value $(-\infty,0]$, but the lhs is restricted to the interval [3/2,3][-3/2,3]. This poses restrictions on the probability distribution $q=(q_{0},q_{1},q_{2})$ in the sense that from the set of all distributions that can generate unistochastic matrices, i.e., those that satisfy the triangle inequalities (13), only distributions with index $\mu\in[1/3,5/9]$ can generate unistochastic matrices from $\mathcal{V}^{\dagger}\mathcal{U}$. The resulting feasible region is depicted in the Fig. 2(b).
Note, however, that although the index condition $\mu\leq 5/9$ is necessarily to generate a bistochastic matrix from the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$, it is not sufficient. Indeed, the conditions (22) to (24) pose some chain-link conditions on four segments $q_{0},q_{1},q_{2}$, and $\sqrt{q_{i}}$, i.e., the following six nontrivial conditions hold
$q_{i}\leq q_{j}+q_{k}+\sqrt{q_{j}},$ (26)
for $(i,j,k)$ taken from $(0,1,2)$. After imposing these constraints, the resulting feasible domain is reduced to the region shown in the Fig. 2(c). For the reason become clear soon, we have to know the region in which $\mu\leq 1/2$. Figure 2(d) illustrates the portion of the feasible region for which $\mu\in[1/3,1/2]$. We are now in a position to characterize a complete set of MEBs for $d=3$.
Proposition 2.
Starting from the standard basis $\{|e_{k}^{(1)}\rangle\}_{k=0}^{2}$, the unitary matrices $\mathcal{U}$, $\mathcal{V}$, and $\mathcal{W}$ generate a complete set of four MEBs. Here, $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$ and $\mathcal{W}=\mathcal{D}_{\beta}\mathcal{U}$, where $\mathcal{D}_{\alpha}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha_{1}},{\mathrm{e}}^{i\alpha_{2}}\}$ and $\mathcal{D}_{\beta}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\beta_{1}},{\mathrm{e}}^{i\beta_{2}}\}$, respectively, and $\mathcal{U}$ is defined by Eq. (8). The pairs $(\alpha_{1},\alpha_{2})$ and $(\beta_{1},\beta_{2})$ are distinct and take their values from the following set
$(0,\theta^{\mu}),\qquad(\theta^{\mu},0),\qquad(\theta^{\mu},\theta^{\mu}),$ (27)
where
$\theta^{\mu}=\cos^{-1}{\Big{[}\frac{-\mu}{1-\mu}\Big{]}},$ (28)
for $\mu\in[1/3,1/2]$. Furthermore, the probabilities $(q_{0},q_{1},q_{2})$ are different permutations of $(q^{\mu}_{+},q^{\mu}_{0},q^{\mu}_{-})$ where
$q^{\mu}_{0}$ $=$ $\frac{1-\mu}{2},$ (29)
$q^{\mu}_{\pm}$ $=$ $\frac{1}{4}\Big{[}(1+\mu)\pm\sqrt{(1+\mu)^{2}-4(1-\mu)^{2}}\Big{]}.$ (30)
Table 1 presents all possible choices for the distribution $q=(q_{0},q_{1},q_{2})$ (first column) versus all possible choices for two pairs $(\alpha_{1},\alpha_{2})$ and $(\beta_{1},\beta_{2})$ (first row). For each chosen probability and each chosen pairs of phases, three equivalent distributions for which three unitary matrices $\mathcal{V}^{\dagger}\mathcal{U}$, $\mathcal{W}^{\dagger}\mathcal{U}$, and $\mathcal{W}^{\dagger}\mathcal{V}$ generate equivalent unistochastic matrices are provided. Remarkably, for a given distribution $q=(q_{0},q_{1},q_{2})$, two other distributions corresponding to cyclic permutations $(123)$ and $(321)$, i.e., $q^{(123)}=(q_{1},q_{2},q_{0})$ and $q^{(321)}=(q_{2},q_{0},q_{1})$, have also contribution to generate equivalent unistochastic matrices $\mathcal{Q}$, $\mathcal{Q}^{(123)}$, and $\mathcal{Q}^{(321)}$.
Table 1: Correspondence between possible choices for the distributions $q=(q_{0},q_{1},q_{2})$ (first column), possible choices for the pairs $(\alpha_{1},\alpha_{2})$ and $(\beta_{1},\beta_{2})$ (first row), and the corresponding three equivalent distributions for which three unitary matrices $\mathcal{V}^{\dagger}\mathcal{U}$, $\mathcal{W}^{\dagger}\mathcal{U}$, and $\mathcal{W}^{\dagger}\mathcal{V}$ generate equivalent unistochastic matrices.
$\begin{array}[]{ll}\alpha=(0,\theta^{\mu})\\ \beta=(\theta^{\mu},0)\end{array}$ $\begin{array}[]{ll}\alpha=(0,\theta^{\mu})\\ \beta=(\theta^{\mu},\theta^{\mu})\end{array}$ $\begin{array}[]{ll}\alpha=(\theta^{\mu},0)\\ \beta=(\theta^{\mu},\theta^{\mu})\end{array}$
$(q^{\mu}_{+},q^{\mu}_{0},q^{\mu}_{-})$ $q^{(123)},q^{(321)} q,q$$q^{(321)},q^{(123)}$ $q^{(123)},q^{(321)},q$
$(q^{\mu}_{+},q^{\mu}_{-},q^{\mu}_{0})$ $q^{(321)},q,q^{(123)}$ $^{(321)},q^{(123)},q q,q$$q^{(123)},q^{(321)}$
$(q^{\mu}_{0},q^{\mu}_{+},q^{\mu}_{-})$ $q^{(123)},q^{(321)},q$ $q^{(123)},q,q^{(321)}$ $q^{(321)},q,q^{(123)}$
$(q^{\mu}_{0},q^{\mu}_{-},q^{\mu}_{+})$ $q^{(123)},q^{(321)},q$ $q^{(123)},q,q^{(321)}$ $q^{(321)},q,q^{(123)}$
$(q^{\mu}_{-},q^{\mu}_{+},q^{\mu}_{0})$ $q^{(321)},q,q^{(123)}$ $^{(321)},q^{(123)},q q,q$$q^{(123)},q^{(321)}$
$(q^{\mu}_{-},q^{\mu}_{0},q^{\mu}_{+})$ $q^{(123)},q^{(321)} q,q$$q^{(321)},q^{(123)}$ $q^{(123)},q^{(321)},q$
II.3.1 Proof of Proposition 2
Proof.
Equation (25) imposes a constraint on the allowed values of the phases $\alpha_{1}$ and $\alpha_{2}$ under which two unitary matrices $\mathcal{U}$ and $\mathcal{V}$ are mutually equi-biased. It depends only on $\mu$, but not on where the probability $q=(q_{0},q_{1},q_{2})$ is located in the feasible region. This makes the constraint invariant under the symmetric group $S_{3}$. A contour plot of this equation in $\alpha_{1}\alpha_{2}$-plane for different values of $\mu$ is illustrated in Fig. 3. The contours which are bounded within the hexagon correspond to $\mu\in[1/3,1/2]$, and those outside the hexagon, i.e., within the two outer triangles, correspond to $\mu\in(1/2,5/9]$.
Figure 3: Contour plot of the lhs of Eq. (25) in $\alpha_{1}\alpha_{2}$-plane for $\alpha_{1},\alpha_{2}\in(-\pi,+\pi)$. (a) The inner boundary (green-ellipse) corresponds to $\mu=1/3$, (b) the outer boundary (red-hexagon) corresponds to $\mu=1/2$, and (c) two little pink contours are plotted for $\mu\approx 5/9$. The exact value for $\mu=5/9$ correspond to two points at $(2\pi/3,-2\pi/3)$ and $(-2\pi/3,2\pi/3)$. The $\alpha_{1}\alpha_{2}$-plane is divided by dashed grid lines to make it easier to identify the quadrant of the angles $\alpha_{1}$ and $\alpha_{2}$.
The feasible region defined above for the probability distribution $q=(q_{0},q_{1},q_{2})$ does not guarantee solutions at any point of the feasible region. Indeed, symmetry conditions requires that the solutions exist only within a measure zero subset of the feasible region, defined by six curves. To see this, note that although the constraint (25) possesses the full symmetry $S_{3}$, a look at Eqs. (22) to (24) reveals that the LHSs of the set is invariant under a subgroup of $S_{3}$, namely, the cyclic group $C_{3}=\{\mathrm{id},(123),(321)\}$. The set, however, will possess the full symmetry under $S_{3}$ if and only if two out of the three terms $\cos{\alpha_{1}}$, $\cos{\alpha_{2}}$, and $\cos{\alpha_{12}}$ become equal. This results in three possible cases as
  1. (a) :
    $\alpha_{1}=0,\quad\quad\alpha_{2}=\theta^{\mu}$,
  2. (b) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=0$,
  3. (c) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=\theta^{\mu}$,
where, from Eq. (25), $\theta^{\mu}$ is obtained as Eq. (28). The index condition $\mu\leq 1/2$ is obtained from $|\cos{\theta^{\mu}}|\leq 1$. Using this in Eqs. (22) to (24), we get for the case (a) the following set of equations
$q_{\sigma(0)}=\frac{2q_{0}q_{1}}{1-\mu},\quad q_{\sigma(2)}=\frac{2q_{1}q_{2}}{1-\mu},\quad q_{\sigma(1)}=\frac{2q_{0}q_{2}}{1-\mu}.$ (31)
Similarly, for cases (b) and (c), the set of Eqs. (22) to (24) reduce to
$q_{\sigma(0)}=\frac{2q_{0}q_{2}}{1-\mu},\quad q_{\sigma(2)}=\frac{2q_{0}q_{1}}{1-\mu},\quad q_{\sigma(1)}=\frac{2q_{1}q_{2}}{1-\mu},$ (32)
and
$q_{\sigma(0)}=\frac{2q_{1}q_{2}}{1-\mu},\quad q_{\sigma(2)}=\frac{2q_{0}q_{2}}{1-\mu},\quad q_{\sigma(1)}=\frac{2q_{0}q_{1}}{1-\mu},$ (33)
respectively. In each case, the corresponding set of equations gives, up to a permutation, the same set of solutions (29) and (30) for the probabilities $(q_{0},q_{1},q_{2})$. Irrespective of which case (a), (b), or (c) is considered, only three out of six possible permutations $\sigma$, namely, those corresponding to the cyclic group $C_{3}=\{\mathrm{id},(123),(321)\}$, are contributed in the solutions.
To complete the characterization of the set of MEBs, we have to construct the third unitary matrix $\mathcal{W}$ as $\mathcal{W}=\mathcal{D}_{\beta}\mathcal{U}$ where $\mathcal{D}_{\beta}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\beta_{1}},{\mathrm{e}}^{i\beta_{2}}\}$. Similar to the arguments given above, we require that the unitaries $\mathcal{W}\mathcal{U}$ and $\mathcal{W}\mathcal{V}$ generate the unistochastic matrix $\mathcal{Q}$, up to a permutation of its rows and columns. The first requirement is achieved if we select the pair $(\beta_{1},\beta_{2})$ from the cases (a), (b), or (c), mentioned above for $(\alpha_{1},\alpha_{2})$. To ensure that $\mathcal{W}\mathcal{V}$ also generates the unistochastic matrix $\mathcal{Q}$, we have to choose the pairs $(\beta_{1},\beta_{2})$ distinct from the pairs $(\alpha_{1},\alpha_{2})$. All possible solutions are listed in Table 1.
Figure 4: The curves represent the allowed values of the distribution $q=(q_{0}=q^{\mu}_{+}$, $q_{1}=\frac{1-\mu}{2}$, $q_{2}=q^{\mu}_{-})$, or any permutation of its probabilities, when $\alpha_{1}=0$ and $\alpha_{2}=\cos^{-1}[-\mu/(1-\mu)]$. Each curve starts from $\mu=1/3$ at the center of the feasible region and ends at $\mu=1/2$ on the boundary of the feasible region. The circles correspond to regions with constant index $\mu$. Corresponding to each $\mu$, only six points, i.e., the distribution $q=(q_{0}=q^{\mu}_{+}$, $q_{1}=\frac{1-\mu}{2}$, $q_{2}=q^{\mu}_{-})$ and its permutations, are allowed for a given $\mu$.
II.3.2 Example: $d=3$ and $\mu=1/2$
For $\mu=1/2$, and by choosing the nonincreasing order for probabilities, i.e., $q_{0}=q_{+}^{\mu}$, $q_{1}=q_{0}^{\mu}$, and $q_{2}=q_{-}^{\mu}$, we find from Eqs. (29) and (30)
$q_{0}=(3+\sqrt{5})/8,\quad q_{1}=1/4,\quad q_{2}=(3-\sqrt{5})/8.$ (34)
Using these in Eqs (9) and (14) to (16), we get $\eta_{0}=0$, $\eta_{1}=\xi_{0}=\xi_{1}=\pi$. Equation (28) gives $\theta^{\mu}=\pi$, which gives for the case (a)
$\alpha_{1}=\beta_{2}=0,\qquad\alpha_{2}=\beta_{1}=\pi.$ (35)
Using these in Eq. (8), one can find $\mathcal{U}$, from which $\mathcal{V}$ and $\mathcal{W}$ can also be obtained as
$\mathcal{U}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}&\sqrt{q_{2}}\\ \sqrt{q_{1}}&-\sqrt{q_{2}}&-\sqrt{q_{0}}\\ \sqrt{q_{2}}&-\sqrt{q_{0}}&\sqrt{q_{1}}\end{pmatrix}$, (36)
$\mathcal{V}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}&\sqrt{q_{2}}\\ \sqrt{q_{1}}&-\sqrt{q_{2}}&-\sqrt{q_{0}}\\ -\sqrt{q_{2}}&\sqrt{q_{0}}&-\sqrt{q_{1}}\end{pmatrix}$, (37)
$\mathcal{W}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}&\sqrt{q_{2}}\\ -\sqrt{q_{1}}&\sqrt{q_{2}}&\sqrt{q_{0}}\\ \sqrt{q_{2}}&-\sqrt{q_{0}}&\sqrt{q_{1}}\end{pmatrix}$. (38)
Obviously, all three unitaries generate the same unistochastic matrix $\mathcal{Q}$. They are also mutually equi-biased as
$\mathcal{V}^{\dagger}\mathcal{U}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}&-\sqrt{q_{2}}\\ \sqrt{q_{1}}&-\sqrt{q_{2}}&\sqrt{q_{0}}\\ -\sqrt{q_{2}}&\sqrt{q_{0}}&\sqrt{q_{1}}\end{pmatrix}$, (39)
$\mathcal{W}^{\dagger}\mathcal{U}=\begin{pmatrix}\sqrt{q_{1}}&\sqrt{q_{2}}&\sqrt{q_{0}}\\ \sqrt{q_{2}}&\sqrt{q_{0}}&-\sqrt{q_{1}}\\ \sqrt{q_{0}}&-\sqrt{q_{1}}&-\sqrt{q_{2}}\end{pmatrix}$, (40)
$\mathcal{W}^{\dagger}\mathcal{V}=\begin{pmatrix}\sqrt{q_{2}}&\sqrt{q_{0}}&\sqrt{q_{1}}\\ \sqrt{q_{0}}&-\sqrt{q_{1}}&\sqrt{q_{2}}\\ \sqrt{q_{1}}&\sqrt{q_{2}}&-\sqrt{q_{0}}\end{pmatrix}$, (41)
where generate three equivalent unistochastic matrices $\mathcal{Q}$, $\mathcal{Q}^{(123)}$, and $\mathcal{Q}^{(321)}$, respectively.
III Entropic uncertainty relation for MEBs
III.1 Inequality for the probabilities of projective measurements in MEBs of a qudit system
[…]
Theorem 3.
[…]
$\sum_{\alpha=1}^{{L}}\sum_{j=0}^{d-1}\left[p_{j}^{(\alpha)}\right]^{2}\leq f_{L}(q)\bigg{[}{\textrm{Tr}}[\varrho^{2}]-\frac{1}{d}\bigg{]}+\frac{L}{d},$ (42)
where
$f_{L}(q)=1+(L-1)\max_{k}|N_{k,d-k}|.$ (43)
Here, $|N_{k,d-k}|$ is given by
$|N_{k,d-k}|=\bigg{[}\mu+2\sum_{m<m^{\prime}}^{d-1}q_{m}q_{m^{\prime}}\cos\Big{(}\frac{2\pi k(m-m^{\prime})}{d}\Big{)}\bigg{]}^{1/2},$ (44)
where $\mu=\sum_{j=0}^{d-1}q_{j}^{2}$ is defined as the index of coincidence of the distribution $q=(q_{0},\cdots,q_{d-1})$, and the maximum is taken over $k\in\{1,\cdots,(d-1)/2\}$ when $d$ is odd, and over $k\in\{1,\cdots,d/2\}$ when $d$ is even. Obviously, $|N_{k,d-k}|=0$ for MUBs for which $\mu=1/d$, thus leading to the results of MolmerPRA2009 .
Proof.
Building on the method presented in Ref. MolmerPRA2009 , we consider the composite system $\mathcal{H}\otimes\mathcal{H}$ and define the following normalized, but not necessarily orthogonal, vectors
$|\phi\rangle$ $=$ $\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|e_{j}^{(1)}\rangle|e_{j}^{(1)}\rangle^{\ast},$ (45)
$|\phi_{k}^{(\alpha)}\rangle$ $=$ $\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}\omega^{kj}|e_{j}^{(\alpha)}\rangle|e_{j}^{(\alpha)}\rangle^{\ast},$ (46)
where $\omega={\mathrm{e}}^{2\pi i/d}$, $k=1,\cdots,d-1$ and $\alpha=1,\cdots,L$. Here, $|e_{j}^{(\alpha)}\rangle^{\ast}$ denotes the complex conjugate of $|e_{j}^{(\alpha)}\rangle$ with respect to a definite basis, say the standard basis $\{|e_{j}^{(1)}\rangle\}_{j=0}^{d-1}$. With this choice, the first basis is real, i.e., $|e_{j}^{(1)}\rangle^{\ast}=|e_{j}^{(1)}\rangle$ for $j=0,\cdots,d-1$, however, for all other bases the expansion coefficients differ by a complex conjugation when they are linear expanded in terms of the first basis. By definition, all these states are maximally entangled states of the two-qudit system $\mathcal{H}\otimes\mathcal{H}$. However, in general, they are neither orthogonal nor complete. More precisely, although the first vector is orthogonal to every vector of the second class, i.e., $\langle\phi|\phi_{k}^{(\alpha)}\rangle=0$, orthogonality fails to hold between vectors of the second class. This can be seen from the overlap or Gram matrix $G$ defined by
$G_{kk^{\prime}}^{(\alpha\alpha^{\prime})}$ $=$ $\langle\phi_{k}^{(\alpha)}|\phi_{k^{\prime}}^{(\alpha^{\prime})}\rangle$
$=$ $N_{kk^{\prime}}+\delta_{\alpha\alpha^{\prime}}(\delta_{kk^{\prime}}-N_{kk^{\prime}}),$
where $N_{kk^{\prime}}=\frac{1}{d}\sum_{j,j^{\prime}=0}^{d-1}\omega^{k^{\prime}j^{\prime}-kj}q_{j\oplus j^{\prime}}$ for $k,k^{\prime}=1,\cdots,d-1$ form the entries of the Gram submatrix $N$ for $\alpha\neq\alpha^{\prime}$, independent of $\alpha$ and $\alpha^{\prime}$. Noteworthy, the matrix $N$ is anti-diagonal as its nonzero entries can be expressed as
$N_{kk^{\prime}}=\begin{cases}\sum_{m=0}^{d-1}\omega^{k^{\prime}m}q_{m}&\mbox{if }k\oplus k^{\prime}=d\\ 0&\mbox{otherwise}.\end{cases}$ (48)
Obviously, $N_{kk^{\prime}}=0$ for $q_{k\oplus k^{\prime}}=1/d$, therefore the vectors defined in Eq. (46) remain mutually orthogonal when the underlying bases are MUBs.
Crucial to our analysis is the fact that the vectors given by Eqs. (45) and (46) are linearly independent, allowing them to span a $D$-dimensional subspace of $\mathcal{H}\otimes\mathcal{H}$ where D=L(d1)+1D=L(d-1)+1. Defining $\{|\chi_{l}\rangle\}_{l=1}^{d^{2}-D}$ as an orthonormal basis for the orthogonal complement of this subspace, and adding these vectors to the previously defined set of vectors we obtain, in general, a nonorthogonal basis for $\mathcal{H}\otimes\mathcal{H}$ as $\{|\phi\rangle,|\phi_{k}^{(\alpha)}\rangle,|\chi_{l}\rangle\}$ where $k=1,\cdots,d-1$ and $\alpha=1,\cdots,L$, and $l=1,\cdots,d^{2}-D$.
Now, for a given qudit quantum state $\varrho$, we define a two-qudit pure state $\varrho\otimes{\mathbb{I}}]|\phi\rangle$, and expand it in terms of the aforementioned basis of $\mathcal{H}\otimes\mathcal{H}$, we get
$[\varrho\otimes{\mathbb{I}}]|\phi\rangle=\frac{1}{d}|\phi\rangle+\sum_{\alpha=1}^{L}\sum_{k=1}^{d-1}a_{k}^{(\alpha)}|\phi_{k}^{(\alpha)}\rangle+|\chi\rangle,$ (49)
where the coefficients $a_{k}^{(\alpha)}$ are given by $a_{k}^{(\alpha)}=[G^{-1}\Omega]_{k}^{(\alpha)}$. A proof can be found in Appendix A. Here, $G^{-1}$ refers to the inverse of the Gram matrix $G$ defined by Eq. (III.1), and the L(d1)L(d-1)-dimensional vector $\Omega=[\Omega_{l}^{(\beta)}]$ is given by
$\Omega_{l}^{(\beta)}=\frac{1}{d}\sum_{j=0}^{d-1}{\omega^{\ast}}^{lj}p_{j}^{(\beta)},$ (50)
for $\beta=1,\cdots,L$ and $l=1,\cdots,d-1$. Based on this, the squared norm of the vector defined by Eq. (49) can be written as
$\langle\phi|[\varrho^{2}\otimes{\mathbb{I}}]|\phi\rangle$ $=$ $\frac{1}{d}{\textrm{Tr}}[\varrho^{2}]$
$=$ $\frac{1}{d^{2}}+\Omega^{\dagger}G^{-1}\Omega+\langle\chi|\chi\rangle.$
From this equation, we get
$\frac{1}{d}\Big{[}{\textrm{Tr}}[\varrho^{2}]-\frac{1}{d}\Big{]}$ $=$ $\Omega^{\dagger}G^{-1}\Omega+\langle\chi|\chi\rangle$
$\geq$ $\Omega^{\dagger}G^{-1}\Omega$
$\geq$ $\|\Omega\|^{2}\lambda_{\min}(G^{-1}),$
where $\lambda_{\min}(G^{-1})$ is the smallest eigenvalue of $G^{-1}$, and the squared norm $\|\Omega\|^{2}=\Omega^{\dagger}\Omega$ can be expressed as
$\Omega^{\dagger}\Omega$ $=$ $\sum_{\alpha=1}^{L}\sum_{k=1}^{d-1}[\Omega_{k}^{(\alpha)}]^{\ast}[\Omega_{k}^{(\alpha)}]$
$=$ $\frac{1}{d^{2}}\sum_{j,j^{\prime}=0}^{d-1}\sum_{\alpha=1}^{L}\sum_{k=1}^{d-1}\omega^{k(j^{\prime}-j)}p_{j^{\prime}}^{(\alpha)}p_{j}^{(\alpha)}$
$=$ $\frac{1}{d^{2}}\sum_{j,j^{\prime}=0}^{d-1}\sum_{\alpha=1}^{L}[d\delta_{jj^{\prime}}-1]p_{j^{\prime}}^{(\alpha)}p_{j}^{(\alpha)}$
$=$ $\frac{1}{d}\sum_{\alpha=1}^{L}\sum_{j=0}^{d-1}[p_{j}^{(\alpha)}]^{2}-\frac{L}{d^{2}}.$
By putting Eq. (III.1) into Eq. (III.1), we get
$\sum_{\alpha=1}^{{L}}\sum_{j=0}^{d-1}\left[p_{j}^{(\alpha)}\right]^{2}$ $\leq$ $\frac{1}{\lambda_{\min}(G^{-1})}\bigg{[}{\textrm{Tr}}[\varrho^{2}]-\frac{1}{d}\bigg{]}+\frac{L}{d}$ (54)
$=$ $\lambda_{\max}(G)\bigg{[}{\textrm{Tr}}[\varrho^{2}]-\frac{1}{d}\bigg{]}+\frac{L}{d},$
where the last line follows from the fact that the Gram matrix of a linearly independent set of vectors is positive definite, as such $\lambda_{\max}(G)=1/\lambda_{\min}(G^{-1})$. To complete the proof, it remains only to determine the eigenvalues of the Gram matrix $G$. A complete derivation of the eigenvalues is provided in Appendix B—from which the maximum eigenvalue we need here is given by $\lambda_{\max}(G)=f_{L}(q)$, where $f_{L}(q)$ is defined by Eq. (43). This concludes the proof. ∎
An important feature that follows from the proof is present here; the Gram matrix $G$, so the submatrix $N$, provides all information on the MEBs, meanwhile, the vector $\Omega\in\mathbb{R}^{d(L-1)}$ contains the probabilities of obtaining different outcomes when the system—initially prepared in state $\varrho$—is measured in MEBs. This feature enables us to parameterize the probability distribution $q=(q_{0},\cdots,q_{d-1})$ associated with each MEBs, using d1d-1 independent parameters provided by the entries of the matrix $N$. See Appendix C.
Quantum incompatibility imposes limitation on the information that can be extracted from measurements of different observables, and the uncertainty relation quantifies how the incompatibility of measurements restricts the precision with which we can simultaneously obtain information about incompatible observables. Entropic uncertainty relation provides a limit on how much information you can gain from measuring incompatible measurements. The inequality given by Theorem 3 for probabilities of projective measurement in MEBs can be used to construct an entropic uncertainty inequality.
Theorem 4.
[…]
$\sum_{\alpha=1}^{{L}}H(\mathcal{B}^{(\alpha)}|\varrho)$ $\geq$ $[L-KC](K+1)\log{(K+1)}$ (55)
- $[L-(K+1)C]K\log{K},$
where $K=\lfloor\frac{L}{C}\rfloor$, and $C$ is the upper bound for $\sum_{\alpha=1}^{{L}}\sum_{j=0}^{d-1}\big{[}p_{j}^{(\alpha)}\big{]}^{2}$. Here for MEBs, $C=f_{L}(q)\big{[}{\textrm{Tr}}[\varrho^{2}]-\frac{1}{d}\big{]}+\frac{L}{d}$ is the upper bound given in Eq. (42).
Proof.
The proof follows exactly the same steps as Ref. MolmerPRA2009 , where the result was established for MUBs based on an inequality by Harremoës and TopsØe HarremosIEEE2001 . Since our proof is identical in structure except for the use of a different upper bound $C$ for $\sum_{\alpha=1}^{{L}}\sum_{j=0}^{d-1}\big{[}p_{j}^{(\alpha)}\big{]}^{2}$, we omit the details here and refer the reader to Ref. MolmerPRA2009 for the complete argument. ∎
IV Entanglement detection
A powerful tool for detecting entanglement is the use of positive but not completely positive maps. A linear map $\Phi:\mathcal{B}(\mathcal{H}_{1})\rightarrow\mathcal{B}(\mathcal{H}_{2})$ is called positive if $\Phi X\geq 0\;\forall X\geq 0$, where $\mathcal{B}(\mathcal{H})$ denotes the space of bounded operators acting on the Hilbert space $\mathcal{H}$. The map $\Phi$ is called completely positive if and only if its trivial extension $\mathrm{id}_{n}\otimes\Phi$ also remains positive for all dimension $n$. For a given positive but not completely positive map $\Phi$, one can construct the corresponding entanglement witness $W=(d-1)(\textrm{id}_{n}\otimes\Phi)|\phi_{d}^{+}\rangle\langle\phi_{d}^{+}|$, where $|\phi_{d}^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d}|ii\rangle$. It turns out that $W$ has nonnegative expectation value on any separable state $\rho$, i.e. ${\textrm{Tr}}{\rho W}\geq 0$ for all separable states. Accordingly, a state for which ${\textrm{Tr}}{\rho W}<0$ is entangled. For a given set of $L$ MUBs, the authors of ChruscinskiPRA2018 have recently introduced the following positive and trace preserving map
$\phi X=\phi_{*}X-\dfrac{1}{d-1}\sum_{\alpha=1}^{{L}}\sum_{k,l=0}^{d-1}\mathcal{O}_{k,l}^{(\alpha)}{\textrm{Tr}}(\tilde{X}P_{l}^{(\alpha)})P_{k}^{(\alpha)},$ (56)
[…]
$W=\dfrac{d+L-1}{d}{\mathbb{I}}\otimes{\mathbb{I}}-\sum_{\alpha=1}^{{L}}\sum_{k,l=0}^{d-1}\mathcal{O}_{kl}^{(\alpha)}\overline{P}_{l}^{(\alpha)}\otimes P_{k}^{(\alpha)},$ (57)
where $\overline{P}_{l}^{(\alpha)}$ is the conjugation of $P_{l}^{(\alpha)}$ and $1\leq L\leq d+1$. The following proposition presents a positive map in terms of a set of $L$ MEBs for a $d$-dimensional system.
Proposition 5.
[…]
$\Phi X$ $=$ $\left[\frac{d+\gamma L-1}{d(d-1)}\right]{\textrm{Tr}}[X]{\mathbb{I}}_{d}$
- $\frac{\gamma}{d-1}\sum_{\alpha=1}^{{L}}\sum_{k,l=0}^{d-1}\mathcal{O}_{k,l}^{(\alpha)}{\textrm{Tr}}[XP_{l}^{(\alpha)}]P_{k}^{(\alpha)},$
is positive and trace preserving. Here, $\gamma$ is given by
$\gamma=\bigg{[}\frac{d-1}{(d-1)f_{L}(q)+L(L-1)(dq_{\max}-1)}\bigg{]}^{1/2},$ (59)
where $q_{\max}=\max_{k}\{q_{0},\cdots,q_{d-1}\}$ and $f_{L}(q)$ is defined by Eq. (43). Also, $\mathcal{O}^{(b)}$ is a set of orthogonal rotation in $\mathbb{R}^{d}$ around the axis $\boldsymbol{n}_{\ast}=(1,1,\cdots,1)/\sqrt{d}$, so that $\mathcal{O}^{(b)}\boldsymbol{n}_{\ast}=\boldsymbol{n}_{\ast}$.
Proof.
First note that the map $\Phi$, defined by Eq. (5), is trace preserving, i.e., ${\textrm{Tr}}[\Phi X]={\textrm{Tr}}[X]$. To show that it is also positive, we require that $\Phi X\geq 0$ for all $X\geq 0$. According to the Mehta’s lemma, a Hermitian matrix $A$ is positive if ${\textrm{Tr}}[A^{2}]\leq{\textrm{Tr}}[A]/(d-1)$. In view of this, the map $\Phi$ is positive if and only if ${\textrm{Tr}}[\Phi\varrho]^{2}\leq{\textrm{Tr}}[\varrho]/(d-1)$ for all rank-one projection operators $P=|\psi\rangle\langle\psi|$. After some tedious but straightforward calculations one finds that
${\textrm{Tr}}[\Phi P]^{2}=\left[\frac{(d-1)^{2}-\gamma^{2}L^{2}}{d(d-1)^{2}}\right]+\frac{\gamma^{2}}{(d-1)^{2}}\textrm{T},$ (60)
where the expression T is defined by
T $=$ $\sum_{\alpha=1}^{{L}}\sum_{l=0}^{d-1}\left\{{\textrm{Tr}}[PP_{l}^{(\alpha)}]\right\}^{2}$
$+$ $\sum_{\alpha\neq\alpha^{\prime}}^{{L}}\sum_{k,l=0}^{d-1}\sum_{k^{\prime},l^{\prime}=0}^{d-1}\mathcal{O}_{k,l}^{(\alpha)}\mathcal{O}_{k^{\prime},l^{\prime}}^{(\alpha^{\prime})}{\textrm{Tr}}[PP_{l}^{(\alpha)}]{\textrm{Tr}}[PP_{l^{\prime}}^{(\alpha^{\prime})}]q_{k\oplus k^{\prime}}.$
Remarkably, for a set of $L$ MUBs, the second term of the above expression reduces to L(L1)/dL(L-1)/d and the first term is bounded from above by 1+(L1)/d1+(L-1)/d, so that the condition ${\textrm{Tr}}[\Phi P]^{2}\leq 1/(d-1)$ is satisfied for $\gamma=1$. For MEBs, on the other hand, the second term is bounded from above by $L(L-1)q_{\max}$ where $q_{\max}=\max_{k}\{q_{k}\}$. For the fist term, however, the upper bound is given by
$\sum_{\alpha=1}^{{L}}\sum_{l=0}^{d-1}\left\{{\textrm{Tr}}[P_{\psi}P_{l}^{(\alpha)}]\right\}^{2}\leq\frac{(d-1)f_{L}(q)+L}{d},$ (62)
where $f_{L}(q)$ is defined by Eq. (43) (see Theorem 3 and its proof). Putting everything together, one can find that the map $\Phi$ is positive if and only if $\gamma$ is given by Eq. (59). This completes the proof that the map $\Phi$ is positive and trace preserving. ∎
Corresponding to the map $\Phi$, defined by Eq. (5), one can define the entanglement witness as follows
$W=\dfrac{d+\gamma L-1}{d}{\mathbb{I}}\otimes{\mathbb{I}}-\gamma\sum_{\alpha=1}^{{L}}\sum_{k,l=0}^{d-1}\mathcal{O}_{kl}^{(\alpha)}\overline{P}_{l}^{(\alpha)}\otimes P_{k}^{(\alpha)}.$ (63)
For illustration, we consider an example of isotropic states of a $d\times d$ system, defined by
$\varrho_{I}=\omega|\phi^{+}_{d}\rangle\langle\phi^{+}_{d}|+\frac{1-\omega}{d^{2}}{\mathbb{I}}_{d}\otimes{\mathbb{I}}_{d},$ (64)
where $0\leq\omega\leq 1$ and $|\phi_{d}^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d}|ii\rangle$. It is shown that isotropic states are entangled if and only if $\omega>1/(d+1)$. Straitforward calculations show that
${\textrm{Tr}}[\varrho_{I}W]$ $=$ $\frac{d+L\gamma\omega-1}{d}-\frac{\gamma\omega}{d}\sum_{\alpha=1}^{L}\sum_{k,l=0}^{d-1}\mathcal{O}_{kl}^{(\alpha)}{\textrm{Tr}}[P_{l}^{(\alpha)}P_{k}^{(\alpha)}]$
$=$ $\frac{d+L\gamma\omega-1}{d}-\frac{\gamma\omega}{d}\sum_{\alpha=1}^{L}{\textrm{Tr}}[\mathcal{O}^{(\alpha)}]$
$=$ $\frac{(d-1)}{d}L\gamma\omega,$
where in the the first equality we have used Eq. (2), and the second equality follows from the fact that the optimum detection happens when ${\textrm{Tr}}[\mathcal{O}^{(\alpha)}]$ takes its maximum value $d$, i.e., when $\mathcal{O}^{(\alpha)}={\mathbb{I}}_{d}$ for all $\alpha$. It turns out that $W$ can detect entanglement of the isotropic states when $\omega>\frac{1}{L\gamma}$. Figure 5 illustrates the behaviour of this function in terms of the index $\mu$ for $d=3$ and different values of $L=2,3,4$. In all cases, the solid curves denote the bounds when MEBs are used to construct witnesses. The dashed lines, on the other hand, captures the bounds when a set of $L$ MUBs are used to construct witnesses. Not surprisingly, for a fixed $L$ the MUBs-based witness is finer than the MEBs-based witness, however, if a MEBs-based witness employs a larger set of measurements, it can surpass the MUBs-based witness in detecting entanglement, particularly within the specific index $\mu$ of the underlying distribution $q$ of the MEBs.
Figure 5: Plots of $(L\gamma)^{-1}$ in terms of index $\mu$ for $d=3$ and different values of $L=2,3,4$. For equal $L$s, the MUBs-based witness (dashed lines) always captures a wider range of states relative to the MEBs-based witness (solid curves). However, if a MEBs-based witness employs a larger set of measurements, it can surpass the MUB-based witness in detecting entanglement, particularly within low values of the index $\mu$.
V Conclusion
In this work, we introduced the concept of mutually equi-biased bases (MEBs), a generalization of the mutually unbiased bases (MUBs). For such bases, we relaxed the condition of mutual maximum unpredictability that holds for MUBs; instead, we associated a probability distribution $q=(q_{0},\cdots,q_{d-1})$ with each of the MEBs. This distribution captures the predictability of the outcomes of a measurement in one basis when the system is prepared in a state from the other basis. However, the information extracted from the measurement is unbiased, meaning that it cannot reveal which state from which basis the system is prepared in, and which basis is used for the measurement.
We derived a complete set of MEBs for $d=2$ and $d=3$, and found that although for $d=2$ a complete set of three MEBs exist for all values of the index $\mu=\sum_{k=0}^{d-1}q_{k}^{2}$, for $d\geq 3$ this is not the case. More precisely, for $d=3$, the constraints imposed by the mutual equi-biasedness limits the range of the index $\mu$ to the feasible region $1/3\leq\mu\leq 1/2$ within which a set of four MEBs exists. Furthermore, in addition to this restriction on $\mu$, for a given $\mu$ in the feasible region the components of the probability distribution $q=(q_{0},q_{1},q_{2})$ are also limited to some specific discrete points; only one probability distribution qq—up to permutation of its components—can generate a complete set of four MEBs.
For a qudit system, we derived an inequality for the probabilities of projective measurements in MEBs. We further used the inequality to derive an entropic uncertainty inequality in MEBs. As expected, the maximal incompatibility inherent to MUBs breaks down in the case of MEBs of the same size. The inequality for probabilities is then used to obtain a class of positive maps, from which we constructed entanglement witnesses.
We examined our entanglement witness for $d\times d$ isotropic states, and, in the case $d=3$, we evaluated it using a set of $L$ MEBs, and presented the results through illustrative plots for $L=2,3,4$. We showed that in the absence of a complete set of MUBs, having a set of MEBs with at least one additional basis can already enhance the construction of finer entanglement witnesses.
These results provide a deeper understanding of the structure and properties of MEBs, offering new tools for entanglement detection and quantum measurement theory. The challenging dimension $d=6$ is currently under investigation to determine whether there exist values of the index $\mu$ for which one can construct a set of at least four MEBs—exceeding a set of three MUBs currently known for this dimension. We hope that our results will have broader applications in quantum information theory; in particular, for dimensions where no more than three MUBs are known, even a set of four MEBs with $\mu>1/d$ could shed light on the challenging open problem of constructing complete set of MUBs.
acknowledgment
This work is based upon research funded by Iran National Science Foundation (INSF) under project No. 4021765.
Appendix A Proof of Eq. (49)
Vectors given by Eq. (46) are linearly independent and span an L(d1)L(d-1)-dimensional subspace of $\mathcal{H}\otimes\mathcal{H}$. Accordingly, corresponding to the nonorthogonal vectors (46), one can define the reciprocal vectors $\{|\widetilde{\phi}_{k}^{(\alpha)}\rangle\}$ such that $\langle\phi_{k}^{(\alpha)}|\widetilde{\phi}_{k^{\prime}}^{(\alpha^{\prime})}\rangle=\delta_{kk^{\prime}}\delta^{(\alpha\alpha^{\prime})}$. The reciprocal vectors can be expanded in terms of the original vectors $|\phi_{k}^{(\alpha)}\rangle$ as
$|\widetilde{\phi}_{k}^{(\alpha)}\rangle=\sum_{k^{\prime}=1}^{d-1}\sum_{\alpha^{\prime}=1}^{L}C_{kk^{\prime}}^{(\alpha\alpha^{\prime})}|\phi_{k^{\prime}}^{(\alpha^{\prime})}\rangle.$ (65)
By multiplying $\langle\phi_{l}^{(\beta)}|$ from the left on this equation, we get
$\delta_{kl}\delta^{(\alpha\beta)}$ $=$ $\sum_{k^{\prime}=1}^{d-1}\sum_{\alpha^{\prime}=1}^{L}C_{kk^{\prime}}^{(\alpha\alpha^{\prime})}G_{lk^{\prime}}^{(\beta\alpha^{\prime})}$
$=$ $\left[CG^{\textrm{T}}\right]_{kl}^{(\alpha\beta)}.$
This implies that $CG^{\textrm{T}}={\mathbb{I}}_{L(d-1)}$, or equivalently, $C^{\textrm{T}}=G^{-1}$, where T stands for the transposition. Accordingly, up to a transposition, the matrix of the coefficients is obtained by the inverse of the Gram matrix $G$. More precisely
$|\widetilde{\phi}_{k}^{(\alpha)}\rangle=\sum_{k^{\prime}=1}^{d-1}\sum_{\alpha^{\prime}=1}^{L}[G^{-1}]_{k^{\prime}k}^{(\alpha^{\prime}\alpha)}|\phi_{k^{\prime}}^{(\alpha^{\prime})}\rangle.$ (67)
Now, for a given qudit quantum state $\varrho$, we define a two-qudit pure state $\varrho\otimes{\mathbb{I}}]|\phi\rangle$, and expand it in terms of the aforementioned basis of $\mathcal{H}\otimes\mathcal{H}$, we get
$[\varrho\otimes{\mathbb{I}}]|\phi\rangle=\frac{1}{d}|\phi\rangle+\sum_{\alpha=1}^{L}\sum_{k=1}^{d-1}a_{k}^{(\alpha)}|\phi_{k}^{(\alpha)}\rangle+|\chi\rangle,$ (68)
where the first term follows from $\langle\phi|[\varrho\otimes{\mathbb{I}}]|\phi\rangle=1/d$, in conjunction with Eq. (45). The unnormalized vector $|\chi\rangle$ represents the component of $\varrho\otimes{\mathbb{I}}]|\phi\rangle$ that lies in the orthogonal subspace spanned by $\{|\chi_{l}\rangle\}$. For the second term, we have
$a_{k}^{(\alpha)}$ $=$ $\langle\widetilde{\phi}_{k}^{(\alpha)}|[\varrho\otimes{\mathbb{I}}]|\phi\rangle$ (69)
$=$ $\frac{1}{d}\sum_{\beta=1}^{L}\sum_{l=1}^{d-1}[G^{-1}]_{kl}^{(\alpha\beta)}\sum_{j=0}^{d-1}{\omega^{\ast}}^{lj}p_{j}^{(\beta)}$
$=$ $[G^{-1}\Omega]_{k}^{(\alpha)},$
[…]
Appendix B Derivation of the eigenvalues of $G$
To proceed with the calculation of eigenvalues of $G$, we first note that the Gram matrix $G$ can be represented in a block form as
$G=\begin{pmatrix}{\mathbb{I}}_{d-1}&N_{d-1}&\cdots&N_{d-1}\\ N_{d-1}&{\mathbb{I}}_{d-1}&\cdots&N_{d-1}\\ \vdots&\vdots&\ddots&\vdots\\ N_{d-1}&N_{d-1}&\cdots&{\mathbb{I}}_{d-1}\end{pmatrix}$, (70)
where ${\mathbb{I}}_{d-1}$ is the unit matrix of dimension (d1)(d-1), and the $(d-1)\times(d-1)$ matrix $N_{d-1}$, defined by Eq. (48), forms the blocks of the Gram matrix $G$. Note here that, to maintain clarity, we explicitly denote the size of each submatrix with a subscript. This simple block form of $G$ allows us to provide the following direct product representation
$G={\mathbb{I}}_{L}\otimes{\mathbb{I}}_{d-1}+S_{L}\otimes N_{d-1},$ (71)
where $S_{L}$ is an $L\times L$ matrix with all off-diagonal entries equal to unity, and the diagonal entries are equal to zero. Both $S_{L}$ and $N_{d-1}$ can be diagonalized unitarily as
$S_{L}=Q_{L}D_{L}Q_{L}^{\dagger},\qquad N_{d-1}=U_{d-1}\Lambda_{d-1}U_{d-1}^{\dagger}.$ (72)
Here
$D_{L}$ $=$ ${\mathrm{diag}}\{L-1,-1,\cdots,-1\},$ (73)
$\Lambda_{d-1}$ $=$ ${\mathrm{diag}}\{\lambda_{1},-\lambda_{1},\cdots,\lambda_{\lfloor\frac{d-1}{2}\rfloor},-\lambda_{\lfloor\frac{d-1}{2}\rfloor},\lambda_{\frac{d}{2}}\},$ (74)
where $\lambda_{i}=|N_{i,d-i}|$, and $\lfloor\frac{d-1}{2}\rfloor$ denotes the integral part of $\frac{d-1}{2}$, i.e., $\lfloor\frac{d-1}{2}\rfloor=\frac{d-1}{2}$ if $d$ is odd, and $\lfloor\frac{d-1}{2}\rfloor=\frac{d}{2}-1$ if $d$ is even. Note that all d1d-1 eigenvalues of $N_{d-1}$ occur in positive-negative pairs, except for the last one, which appears alone. Obviously, this lone eigenvalue, i.e., $\lambda_{\frac{d}{2}}$, exists only when $d$ is even.
Having these diagonal forms, one can write $G$ as
$G=[Q_{L}\otimes U_{d-1}]G_{\textrm{diag}}[Q_{L}^{\dagger}\otimes U_{d-1}^{\dagger}],$ (75)
where $G_{\textrm{diag}}$ is the diagonal representation of the Gram matrix $G$,
$G_{\textrm{diag}}={\mathbb{I}}_{L}\otimes{\mathbb{I}}_{d-1}+D_{L}\otimes\Lambda_{d-1}.$ (76)
To classify the eigenvalues of $G$, we discuss even and odd cases, separately.
Eigenvalues of $G$ when $d$ is odd.—In this case, for $i\in\{1,\cdots,(d-1)/2\}$, the nondegenerate eigenvalues are
$1\pm(L-1)|N_{i,d-i}|,$ (77)
while the degenerate ones, each with multiplicity L1L-1, are
$1\pm|N_{i,d-i}|.$ (78)
Eigenvalues of $G$ when $d$ is even.—In this case, the eigenvalues are the same as above, here for $i\in\{1,\cdots,d/2-1\}$, along with the two new eigenvalues $1+(L-1)|N_{d/2,d/2}|$ and $1-|N_{d/2,d/2}|$, which have multiplicities $1$ and L1L-1, respectively.
Now, let $\lambda_{\max}(G)$ be the maximum eigenvalue of the Gram matrix $G$. For both cases of even and odd $d$, we find
$\lambda_{\max}(G)=1+(L-1)\max_{i}|N_{i,d-i}|.$ (79)
Appendix C Parametrization of $q=(q_{0},\cdots,q_{d-1})$ for MEBs
For any dimension $d$, the nonzero entries of $N$ occurs when $k\oplus k^{\prime}=d$, as follows
$N_{kk^{\prime}}$ $=$ $\sum_{m=0}^{d-1}\omega^{k^{\prime}m}q_{m}=\sum_{m=0}^{d-1}\omega^{-km}q_{m}$
$=$ $\delta_{k}^{\textrm{R}}+i\delta_{k}^{\textrm{I}},$
where $\delta_{k}^{\textrm{R}}$ and $\delta_{k}^{\textrm{I}}$ are defined as the real and imaginary parts of $N_{kk^{\prime}}$, respectively, as
$\delta_{k}^{\textrm{R}}$ $=$ $\sum_{m=0}^{d-1}\cos\bigg{[}\frac{2\pi km}{d}\bigg{]}q_{m},$ (81)
$\delta_{k}^{\textrm{I}}$ $=$ $-\sum_{m=0}^{d-1}\sin\bigg{[}\frac{2\pi km}{d}\bigg{]}q_{m}.$ (82)
Accordingly, for the absolute values we have
$|N_{kk^{\prime}}|^{2}$ $=$ $[\delta_{k}^{\textrm{R}}]^{2}+[\delta_{k}^{\textrm{I}}]^{2}$
$=$ $\mu+2\sum_{m<m^{\prime}}^{d-1}\cos\bigg{[}\frac{2\pi k(m-m^{\prime})}{d}\bigg{]}q_{m}q_{m^{\prime}}.$
To illustrate the nature of this parametrization, we explicitly consider the cases $d=2,3$, and $4$, as follows.
d=2d=2.—In this case the only nonzero entry of $N$ is given by $N_{11}=\delta$, where $\delta=q_{0}-q_{1}$, such that $1\leq\delta\leq 1$. It follows therefore that
$q_{0}$ $=$ $(1+\delta)/2,$ (84)
$q_{1}$ $=$ $(1-\delta)/2.$
d=3d=3.—For $d=3$, the nonzero entries of $N$ are given by $N_{12}=N_{21}^{\ast}=\delta_{1}+i\delta_{2}$. Two independent parameters are given $\delta_{1}=(3q_{0}-1)/2$ and $\delta_{2}=\sqrt{3}(q_{2}-q_{1})/2$ where we have used them in Sec. II to plot the feasible region in $\delta_{2}\delta_{1}$-plane for $d=3$. Here
$-1/2\leq$ $\delta_{1}$ $\leq 1,$ (85)
$(-1+\delta_{1})/\sqrt{3}\leq$ $\delta_{2}$ $\leq(1-\delta_{1})/\sqrt{3}.$ (86)
We therefore have
$q_{0}$ $=$ $1/3+2\delta_{1}/3,$ (87)
$q_{1}$ $=$ $1/3-\delta_{1}/3-\delta_{2}/\sqrt{3},$
$q_{2}$ $=$ $1/3-\delta_{1}/3+\delta_{2}/\sqrt{3}.$
In this parametrization, one can write
$|N_{1,2}|=\sqrt{\delta_{1}^{2}+\delta_{2}^{2}}=\sqrt{(3\mu-1)/2}.$ (88)
d=4d=4.—When $d=4$, the nonzero entries of matrix $N$ are given by $N_{13}=N_{31}^{\ast}=\delta_{1}+i\delta_{2}$ and $N_{22}=\delta_{3}$. Therefore, three independent parameters are given by $\delta_{1}=q_{0}-q_{2}$, $\delta_{2}=q_{3}-q_{1}$, and $\delta_{3}=q_{0}-q_{1}+q_{2}-q_{3}$, such that
$-(1+\delta_{3})/2\leq$ $\delta_{1}$ $\leq(1+\delta_{3})/2,$ (89)
$-(1-\delta_{3})/2\leq$ $\delta_{2}$ $\leq(1-\delta_{3})/2,$ (90)
$-1\leq$ $\delta_{3}$ $\leq 1.$ (91)
It follows therefore that
$q_{0}$ $=$ $1/4+\delta_{1}/2+\delta_{3}/4,$
$q_{1}$ $=$ $1/4-\delta_{2}/2-\delta_{3}/4,$
$q_{2}$ $=$ $1/4-\delta_{1}/2+\delta_{3}/4,$
$q_{3}$ $=$ $1/4+\delta_{2}/2-\delta_{3}/4.$